# Enhancing Fairness through Reweighting: A Path to Attain the Sufficiency Rule

### v1

@@ 1 Introduction
In practical terms, neglecting the sufficiency rule can result in significant biases within intelligent healthcare systems. For instance, many health systems utilize algorithms to identify and support patients with complex health requirements. These algorithms generate a score indicating the level of healthcare needs, with higher scores suggesting greater sickness and the need for more care. Notably, a study by [32] uncovers a widely used industry algorithm affecting millions of patients, which exhibits pronounced racial bias. It was found that for a given predicted score $\hat{Y}=s$, black patients tend to be considerably sicker than white patients ($\mathbb{E}_{black}[Y|\hat{Y}=s]>\mathbb{E}_{white}[Y|\hat{Y}=s]$). Moreover, the study highlights that rectifying this disparity could significantly increase the percentage of black patients receiving additional care from 17.7% to 46.5%. From an algorithmic perspective, the sufficiency rule is generally incompatible with concepts such as independence or separation, as demonstrated in Section 2 and the Appendix. This suggests that existing fair algorithms designed for independence or separation may not enhance or could even exacerbate issues related to the sufficiency rule. Notably, recent work on Invariant Risk Minimization (IRM) proposed by [2, 4] has potential to address this challenge. IRM seeks to maintain invariant correlations between the embedding (or representation) and the true label by incorporating regularization techniques into Deep Neural Network (DNN) training. The criteria of the sufficiency rule and IRM are intrinsically consistent (see more details in Section 3.1.1); the idea being that if the correlations between the embedding (or representation) and the true label remain robust and unaffected by specific sub-groups, the resulting representation can be considered fair. To better understand this concept, IRM addresses the challenge of ensuring that a cow is correctly classified as a cow in a picture, regardless of whether the background is Grass or Desert [2]. On the other hand, the sufficiency rule aims to ensure that a patient predicted to be high-risk is truly high-risk, regardless of whether the patient is Black or White. IRM approaches have attracted attention due to their promising performance on modest models and datasets [2] and their simplicity in facilitating end-to-end training. Nevertheless, recent studies have indicated diminished effectiveness of the regularization terms when applied to overparameterized DNNs [9, 25]. For instance, CelebA comprises only 200k training data, whereas ResNet-18 boasts 11.4 million parameters. Overparameterized DNNs can easily diminish the regularization term of IRM to zero during training while still depending on spurious features. In such scenarios, applying IRM methods directly for fairness to uphold the sufficiency rule in relatively larger models is deemed inappropriate.
@@ 2.1 Sufficiency Rule in Fairness
$\mathbb{E}_{\mathcal{D}_{0}}[Y|\hat{Y}=s]=\mathbb{E}_{\mathcal{D}_{1}}[Y|\hat{Y}=s],\forall s\in\mathcal{Y}$ (1)
@@ 2.1.2 Relation to Other Fairness Notions
We briefly contrast the Sufficiency rule with the commonly employed Independence and Separation rules in binary classification. For comprehensive justifications and comparisons, please consult Appendix.
@@
$\mathbb{E}_{\mathcal{D}_{0}}[\hat{Y}|Y=s]=\mathbb{E}_{\mathcal{D}_{1}}[\hat{Y}|Y=s],\forall t\in Y$ (4)
In binary classification, the Separation rule is also referred to as Equalized Odds (EO) [18]. Additionally, [3] have further illustrated that if $P_{\mathcal{D}_{0}}(Y=y)\neq P_{\mathcal{D}_{1}}(Y=y)$ and the joint distribution of $(Y,\hat{Y})$ has a positive probability in $\mathcal{D}_{0}$ and $\mathcal{D}_{1}$, then it is impossible for both the Sufficiency and Separation rules to coexist [5] (please refer to Appendix for further details).
@@ 2.2 Invariant Risk Minimization
IRM operates under the assumption that there are multiple environments :=$\mathcal{E}:=\{e_{1},e_{2},...,e_{E}\}$ within the sample space $\mathcal{X}\times\mathcal{Y}$, each characterized by distinct joint distributions. Furthermore, it assumes that the correlation between the spurious features and labels varies inconsistently across these environments. The predictor $f(\cdot;\theta)$ in IRM is expressed as a composite function of a representation $\phi(\cdot;\Phi)$ and a classifier $h(\cdot;v)$, formulated as $f(\cdot;\theta)=h(\phi(\cdot;\Phi);v)$, where $\theta=\{v,\Phi\}$ represents the trainable parameters. The fundamental idea is that if a predictor $f(\cdot;\theta)$ performs effectively across all environments, it suggests that the correlation between the spurious features and labels is not accurately captured[35, 2]. In these cases, the data representation function $\phi$ elicits an invariant predictor across environments $\mathcal{E}$ if and only if for all latent $z$ in the intersection of the supports of $\phi(X^{e})$ we have $\mathbb{E}[Y^{e}|\phi(X^{e})=z]=\mathbb{E}[Y^{e^{\prime}}|\phi(X^{e^{\prime}})=z]$, for all $e,e^{\prime}\in\mathcal{E}$ (For loss functions such as the mean squared error and the cross-entropy, optimal classifiers can be written as conditional expectations). Please refer to Appendix for more details. Consequently, IRM aims to minimize a specific IRM risk to identify such a robust predictor. Several approaches have been proposed to enhance IRM: [23, 40] advocate for penalizing the variance of risks across different environments, while [7, 41] attempt to estimate the violation of invariance by training neural networks. Moreover, theoretical guarantees for IRM on linear models with adequate training environments are provided by [2, 36, 8].

### v2

@@ 1 Introduction
In practical terms, neglecting the sufficiency rule can result in significant biases within intelligent healthcare systems. For instance, many health systems utilize algorithms to identify and support patients with complex health requirements. These algorithms generate a score indicating the level of healthcare needs, with higher scores suggesting greater sickness and the need for more care. Notably, a study by [32] uncovers a widely used industry algorithm affecting millions of patients, which exhibits pronounced racial bias. It was found that for a given predicted score $\hat{Y}=t$, black patients tend to be considerably sicker than white patients ($\mathbb{E}_{black}[Y|\hat{Y}=t]>\mathbb{E}_{white}[Y|\hat{Y}=t]$). Moreover, the study highlights that rectifying this disparity could significantly increase the percentage of black patients receiving additional care from 17.7% to 46.5%. From an algorithmic perspective, the sufficiency rule is generally incompatible with concepts such as independence or separation, as demonstrated in Section 2 and the Appendix [45]. This suggests that existing fair algorithms designed for independence or separation may not enhance or could even exacerbate issues related to the sufficiency rule. Notably, recent work on Invariant Risk Minimization (IRM) proposed by [2, 4] has potential to address this challenge. IRM seeks to maintain invariant correlations between the embedding (or representation) and the true label by incorporating regularization techniques into Deep Neural Network (DNN) training. The criteria of the sufficiency rule and IRM are intrinsically consistent (see more details in Section 3.1.1); the idea being that if the correlations between the embedding (or representation) and the true label remain robust and unaffected by specific sub-groups, the resulting representation can be considered fair. To better understand this concept, IRM addresses the challenge of ensuring that a cow is correctly classified as a cow in a picture, regardless of whether the background is Grass or Desert [2]. On the other hand, the sufficiency rule aims to ensure that a patient predicted to be high-risk is truly high-risk, regardless of whether the patient is Black or White. IRM approaches have attracted attention due to their promising performance on modest models and datasets [2] and their simplicity in facilitating end-to-end training. Nevertheless, recent studies have indicated diminished effectiveness of the regularization terms when applied to overparameterized DNNs [9, 25]. For instance, CelebA comprises only 200k training data, whereas ResNet-18 boasts 11.4 million parameters. Overparameterized DNNs can easily diminish the regularization term of IRM to zero during training while still depending on spurious features. In such scenarios, applying IRM methods directly for fairness to uphold the sufficiency rule in relatively larger models is deemed inappropriate.
@@ 2.1 Sufficiency Rule in Fairness
$\mathbb{E}_{\mathcal{D}_{0}}[Y|\hat{Y}=t]=\mathbb{E}_{\mathcal{D}_{1}}[Y|\hat{Y}=t],\forall t\in\mathcal{Y}$ (1)
@@ 2.1.2 Relation to Other Fairness Notions
We briefly contrast the Sufficiency rule with the commonly employed Independence and Separation rules in binary classification. For comprehensive justifications and comparisons, please consult Appendix [45].
@@
$\mathbb{E}_{\mathcal{D}_{0}}[\hat{Y}|Y=t]=\mathbb{E}_{\mathcal{D}_{1}}[\hat{Y}|Y=t],\forall t\in Y$ (4)
In binary classification, the Separation rule is also referred to as Equalized Odds (EO) [18]. Additionally, [3] have further illustrated that if $P_{\mathcal{D}_{0}}(Y=y)\neq P_{\mathcal{D}_{1}}(Y=y)$ and the joint distribution of $(Y,\hat{Y})$ has a positive probability in $\mathcal{D}_{0}$ and $\mathcal{D}_{1}$, then it is impossible for both the Sufficiency and Separation rules to coexist [5] (please refer to Appendix[45] for further details).
@@ 2.2 Invariant Risk Minimization
IRM operates under the assumption that there are multiple environments :=$\mathcal{E}:=\{e_{1},e_{2},...,e_{E}\}$ within the sample space $\mathcal{X}\times\mathcal{Y}$, each characterized by distinct joint distributions. Furthermore, it assumes that the correlation between the spurious features and labels varies inconsistently across these environments. The predictor $f(\cdot;\theta)$ in IRM is expressed as a composite function of a representation $\phi(\cdot;\Phi)$ and a classifier $h(\cdot;v)$, formulated as $f(\cdot;\theta)=h(\phi(\cdot;\Phi);v)$, where $\theta=\{v,\Phi\}$ represents the trainable parameters. The fundamental idea is that if a predictor $f(\cdot;\theta)$ performs effectively across all environments, it suggests that the correlation between the spurious features and labels is not accurately captured[35, 2]. In these cases, the data representation function $\phi$ elicits an invariant predictor across environments $\mathcal{E}$ if and only if for all latent $z$ in the intersection of the supports of $\phi(X^{e})$ we have $\mathbb{E}[Y^{e}|\phi(X^{e})=z]=\mathbb{E}[Y^{e^{\prime}}|\phi(X^{e^{\prime}})=z]$, for all $e,e^{\prime}\in\mathcal{E}$ (For loss functions such as the mean squared error and the cross-entropy, optimal classifiers can be written as conditional expectations). Please refer to Appendix[45] for more details. Consequently, IRM aims to minimize a specific IRM risk to identify such a robust predictor. Several approaches have been proposed to enhance IRM: [23, 40] advocate for penalizing the variance of risks across different environments, while [7, 41] attempt to estimate the violation of invariance by training neural networks. Moreover, theoretical guarantees for IRM on linear models with adequate training environments are provided by [2, 36, 8].
@@ 5 Discussion and Conclusion
[…]
This work has received funding from the European Union’s Horizon 2020 research and innovation programme under Marie Sklodowska-Curie Actions (grant agreement number 860630) for the project “NoBIAS - Artificial Intelligence without Bias”.